OpenStudy (anonymous):

Solve: -4(x + 1) - 3 = -3(x - 4) -19 19 2 0

OpenStudy (anonymous):

-4x - 4 - 3 = -3x +12 -4x - 7 = -3x +12 -4x +3x = 12+ 7 -x = 19 x = -19